given a model with the following structural quations
y = c + I + G + (X - M) where
C = 30 + 0.75Y
I = N80 million
G = 120 million
X = 20 million
M = 5 + 0.2 y
determine
i) equilibrium national income
ii) income multiplier
b analyze the import demand equation
use a numerical example to define and explain the concept of multipier
1
Expert's answer
2020-12-28T16:35:49-0500
"Y=30+0.75Y+80+120+20-5-0.2Y"

"0.45Y=245"

i)


"Y=544"

ii)


"\mu=0.45"

b)


"M=5+0.2Y"


"M=5+0.2\times544=113.8"